OpenStudy (anonymous):

The range and domain of sqrt 9 - x^2 ?

OpenStudy (anonymous):

u can do it...

OpenStudy (anonymous):

i can find the domain but range!!!

OpenStudy (anonymous):

y = f(x) y^2 = 9-x^2

OpenStudy (anonymous):

@mukushla help

OpenStudy (anonymous):

what is domain?

OpenStudy (anonymous):

range is the y coordinate

OpenStudy (anonymous):

[-3 , 3]

OpenStudy (anonymous):

the domain should be all real nos.

OpenStudy (anonymous):

ok see this \[-3\le x \le 3 \\0\le x^2 \le 9 \\ -9 \le -x^2 \le 0 \\ 0 \le 9-x^2 \le 9 \\ 0 \le y^2 \le 9 \\ 0 \le y \le 3\]

OpenStudy (anonymous):

is taht your query?

OpenStudy (anonymous):

hemicircle...

OpenStudy (anonymous):

@mukushla u r correct but hw???????????

OpenStudy (anonymous):

no @shruti my ques is whole root

OpenStudy (anonymous):

@Romero